OpenStudy (sambam):

Divide the following polynomial 28y^6/14y^4

OpenStudy (anonymous):

\[\huge\frac{\cancel{28}^2}{\cancel{14}^1}\frac{y^{\cancel{6}2}}{y^{\cancel{4}0}}\] \[\huge \frac{2}{1}\frac{y^2}{1}=2y^2\]

OpenStudy (sambam):

thanks so much